# EGG-SR: Embedding Symbolic Equivalence into Symbolic Regression via Equality Graph

### v1

@@ D.1 Implementation of Egg-MCTS
Connection to Existing Research. In many applications, multiple action sequences may lead to the same state, resulting in duplicate nodes within the search tree. To mitigate this redundancy, DBLP:journals/kbs/SaffidineCM12 proposed the use of a transposition table in the context of Go, enabling the reuse of information from previous updates. Similarly, DBLP:conf/acml/LeurentM20 introduced Monte Carlo Graph Search (MCGS), which replaces the tree structure with a graph that merges identical states. In MCGS, nodes correspond to unique states $s\in\mathcal{S}$ and edges represent state transitions, with the root node corresponding to the initial state $s_{0}$. At iteration $n$, if executing action $a$ from state $s$ leads to a successor state $s^{\prime}$ that already exists in the graph, no new node is created.
@@
Our Egg-MCTS adopts the same principle of detecting identical states during search but avoids explicit graph construction. Instead, it simultaneously updates all paths in the tree that lead to the selected state. This design preserves compatibility with standard MCTS implementations while achieving a theoretical acceleration comparable to that of DBLP:conf/acml/LeurentM20.
D.2 Implementation of Egg-DRL
@@
Expression generation begins with the decoder sampling a sequence of grammar rules in an autoregressive manner. The sequential decoder can be implemented using various architectures, such as RNNs (salehinejad2017recent), GRUs (chung2014empirical), LSTMs (greff2016lstm), or decoder-only Transformers (vaswani2017attention). The input and output vocabularies for the decoder consist of grammar rules that encode variables, coefficients, and mathematical operators. Figure 7(a) illustrates an example of the output vocabulary.
@@
At each time step, the model predicts a categorical distribution over the next rule, conditioned on the previously generated tokens. At the $t$-th step, the decoder (denoted $\mathtt{SequentialDecoder}$) takes the previously generated rules $(\tau_{1},\dots,\tau_{t})$ and the hidden state $\mathbf{h}_{t}$, and computes:
$\mathbf{z}_{t}$ $=\mathtt{SequentialDecoder}(\tau_{t},\mathbf{h}_{t}),$
$p_{\theta}(\tau_{t+1}\mid\tau_{1},\dots,\tau_{t})$ $=\mathtt{softmax}(\mathbf{z}_{t}W_{o}+b_{o}),$
@@
where $W_{o}\in\mathbb{R}^{d\times|V|}$ is the output weight matrix, $b_{o}\in\mathbb{R}^{|V|}$ is the bias vector, and $|V|$ is the size of the vocabulary. The next rule $\tau_{t+1}$ is then sampled from this distribution: $\tau_{t+1}\sim p(\tau_{t+1}\mid\tau_{1},\dots,\tau_{t}).$ The sampled rule is fed into the decoder as input for the next step, until the sequence is completed. After $L$ steps, the full sequence $\tau=(\tau_{1},\ldots,\tau_{L})$ is generated with probability $p_{\theta}(\tau)=\prod_{t=1}^{L-1}p_{\theta}(\tau_{t+1}\mid\tau_{1},\dots,\tau_{t})$.
@@
Connection to Existing Research. Several techniques have been introduced to reduce the variance of policy gradient estimates. One widely used approach is the control variate method, where a baseline is subtracted from the reward to stabilize the gradient (DBLP:conf/uai/WeaverT01). Another idea is Rao-Blackwellization (8a19a051-2be1-32f8-8b83-075ff1f85473). Other approaches, such as reward reshaping (DBLP:conf/nips/ZhengOS18), modify rewards for specific state-action pairs. Inspired by stochastic variance-reduced gradient methods (johnson2013accelerating; DBLP:conf/iclr/0002FKLL21), (DBLP:conf/icml/PapiniBCPR18) proposed a variance-reduction technique tailored for policy gradients.
@@
D.3 Implementation of Egg-LLM
Our Egg is adopted on top of the original LLM-SR (DBLP:journals/iclr/llmsr25).
@@ D.3 Implementation of Egg-LLM
    Inferring Equivalent Expressions. For ease of integration with the e-graph system, we convert the generated Python functions into lambda expressions. This simplifies their manipulation and evaluation during equivalence analysis.
    # Original function format
    def hypothesis(x1, x2, coeffs):
    expr = x1 * coeffs[0] + x2 * coeffs[0]
    return expr
[…]
    hypothesis = lambda x1, x2, coeffs: x1 * coeffs[0] + x2 * coeffs[0]
    This conversion allows us to transform the hypothesized Python function into an expression tree directly. For each fitted expression, we construct an e-graph and apply a predefined set of rewrite rules until reaching a fixed iteration limit. From the resulting e-graph, we sample $K$ unique equivalent expressions. Each sampled expression is then converted back into a Python function to facilitate further interaction with the LLM.
    # obtained equivalent expression
    equiva_seq = ["A->log(A)", "A->A*A", "A->x1", "A->x2"]
    # Converted function format
    def equiva_hypothesis(x1, x2, coeffs):
    return x1 * x2
  4. (d)
    Egg-based Feedback. In subsequent iterations, the LLM receives feedback consisting of previously generated expressions along with their fitness scores. Our feedback is enriched with both the original hypotheses and their equivalent expressions derived via Egg module, enabling the model to refine future generations. High-performing expressions are retained and further updated over multiple rounds.
@@
D.4 Final Remark
The E2E-Transformer framework (DBLP:conf/nips/KamiennydLC22) is structurally very similar to the neural network used in the DRL model, with the main difference being that it is trained using a cross-entropy loss function. A straightforward way to integrate the Egg module into E2E-Transformer is through data augmentation, where additional training examples are generated using Egg.
To date, only one work has explored the use of a text-based diffusion model for symbolic regression (bastiani2025diffusion). However, its implementation has not been released publicly. Moreover, the proposed pipeline relies on multiple interconnected components, which makes it difficult to reproduce or isolate for independent evaluation. For these reasons, we do not include a comparison between the text-based diffusion approach and Egg in this paper.
Several extensions of the baseline considered in this study offer empirical improvements but lack theoretical guarantees. For example, DBLP:conf/nips/MundhenkLGSFP21 employ genetic programming to refine the predictions of the DRL model, and DBLP:journals/corr/abs-2303-03192 introduce unit constraints in physical systems to eliminate infeasible expressions. Integrating Egg with these variants is an interesting direction that we leave for future investigation.
Appendix E Experiment Settings
E.1 Baselines and Hyper-parameters Configurations
@@
MCTS. (DBLP:conf/iclr/Sun0W023) propose to use the Monte Carlo tree search algorithm to explore the space of symbolic expressions defined by a context-free grammar (described in Section A) to find high-performing expressions. Despite being implemented similarly, this method appears under various names in previous works (DBLP:conf/icml/TodorovskiD97; DBLP:conf/dis/GanzertGSK10; DBLP:journals/kbs/BrenceTD21; DBLP:conf/iclr/Sun0W023; DBLP:conf/icml/KamiennyLLV23). We choose the implementation  [^5] https://github.com/isds-neu/SymbolicPhysicsLearner from DBLP:conf/iclr/Sun0W023 and serve as a representative of this family of methods.
DRL. (DBLP:conf/iclr/PetersenLMSKK21) propose to use a recurrent neural network (RNN) trained via a (risk-seeking) reinforcement learning objective. The RNN sequentially generates candidate expressions and is optimized using a risk-seeking policy gradient to encourage the discovery of high-quality expressions. We chose this implementation  [^6] https://github.com/dso-org/deep-symbolic-optimization. We use three types of sequential decoders for the time benchmark setting. The major configurations are listed in Table 4.
@@
LLM-SR. (DBLP:journals/iclr/llmsr25) leverage pretrained large language models, such as GPT4o, to generate symbolic expressions based on task-specific prompts. The expressions are evaluated and iteratively refined over multiple rounds. We choose their implementation at  [^7] https://github.com/deep-symbolic-mathematics/LLM-SR.
@@
Expression-related Configurations
When fitting the values of open constants in each expression, we sample a batch of data with batch size $1024$ from the data Oracle. The open constants in the expressions are fitted on the data using the BFGS optimizer [^8] https://docs.scipy.org/doc/scipy/reference/optimize.minimize-bfgs.html. We use a multi-processor library to fit multiple expressions using 8 CPU cores in parallel. This greatly reduced the total training time.
@@
The normalized mean-squared error metric is further defined in Equation 5.
@@ E.2 Evaluation Metrics
where the empirical variance $\sigma_{y}=\sqrt{\frac{1}{n}\sum_{i=1}^{n}\left(y_{i}-\frac{1}{n}\sum_{i=1}^{n}y_{i}\right)^{2}}$. Note that the coefficient of determination ($R^{2}$) metric (nagelkerke1991note; DBLP:conf/nips/CavaOBFVJKM21) is equal to $(1-\text{NMSE})$ and therefore omitted in the experiments.
Appendix F Extra Experimental Results
F.1 Visualization of Egg Construction
@@ I.44.4 $c_{1}x_{0}x_{1}\log(x_{2}/x_{3})$ $c_{1}x_{0}x_{1}(\log(x_{2})-\log(x_{3}))$ Figure 17
F.2 Additional Visualization for Selected Expressions in Feynman Dataset

### v2

@@ B.4.1 Implementation of Egg-MCTS
Connection to equivalence-aware learning in MCTS. In many applications, multiple action sequences may lead to the same state, resulting in duplicate nodes within the search tree. To mitigate this redundancy, DBLP:journals/kbs/SaffidineCM12 proposed the use of a transposition table in the context of Go, enabling the reuse of information from previous updates. Similarly, DBLP:conf/acml/LeurentM20 introduced Monte Carlo Graph Search (MCGS), which replaces the search tree with a graph that merges identical states. In MCGS, nodes correspond to unique states $s\in\mathcal{S}$ and edges represent state transitions, with the root node corresponding to the initial state $s_{0}$. At iteration $n$, if executing action $a$ from state $s$ leads to a successor state $s^{\prime}$ that already exists in the graph, no new node is created.
@@
Our Egg-MCTS adopts the same principle of detecting identical states during search, but avoids explicit graph construction. The constructed e-graph serves as a transposition table that stores and detects identical states. Following the same update rule used in a transposition table, Egg-MCTS simultaneously updates all paths in the tree that lead to the selected state. This design preserves compatibility with standard MCTS implementations while achieving a theoretical acceleration comparable to that of DBLP:conf/acml/LeurentM20.
B.4.2 Implementation of Egg-DRL
Markov Decision Process Formulation for DRL. We model this problem as an RL agent that searches for the optimal sequence of grammar rules. At decoding step $t$, the agent samples a rule $\tau_{t}$ conditioned on the previously selected rules $\tau_{1},\ldots,\tau_{t-1}$. We define the state at step $t$ as the prefix of sampled rules $s_{t}\coloneqq(\tau_{1},\ldots,\tau_{t-1})$, and the grammar rules in the output vocabulary constitute the action space for the RL agent. The sequential decoder, parameterized by $\theta$, defines a stochastic policy
$\pi_{\theta}(a_{t}=a^{\prime}|s_{t})\coloneqq p_{\theta}(\tau_{t}=a^{\prime}|\tau_{1},\ldots,\tau_{t-1}).$
The environment transition is deterministic: applying action $\tau_{t}$ in state $s_{t}$ yields the next state $s_{t+1}=(\tau_{1},\ldots,\tau_{t})$. Rewards are obtained by evaluating the completed expression on the dataset according to a chosen evaluation metric (e.g., $1/(1+\texttt{NMSE}(\phi))$). The MDP has a finite horizon: an episode terminates once the output sequence corresponds to a valid expression with no non-terminal symbols, or when the maximum sequence length $H$ is reached. The objective of the RL agent is to learn a policy that selects sequences of grammar rules so as to maximize the expected reward.
@@ B.4.2 Implementation of Egg-DRL
Configuration of Sequential Decoder. The sequential decoder can be implemented using various architectures, such as RNNs [salehinejad2017recent], GRUs [chung2014empirical], LSTMs [greff2016lstm], or decoder-only Transformers [vaswani2017attention]. The input and output vocabularies for the decoder consist of grammar rules that encode variables, coefficients, and mathematical operators. Figure 8(a) illustrates an example of the output vocabulary.
@@
At each time step, the model predicts a categorical distribution over the next rule, conditioned on the previously generated rules. At step $t$, the decoder (denoted as $\mathtt{SequentialDecoder}$”) takes the previously generated rules $(\tau_{1},\dots,\tau_{t-1})$ and the hidden state $\mathbf{h}_{t-1}$, and computes
$\mathbf{h}_{t}$ $=\mathtt{SequentialDecoder}(\tau_{t-1},\mathbf{h}_{t-1}),$
$\mathbf{z}_{t}$ $=W_{o}\mathbf{h}_{t}+b_{o},$
$p_{\theta}(\tau_{t}|\tau_{1},\dots,\tau_{t-1})$ $=\mathtt{softmax}(\mathbf{z}_{t}),$
@@
where $W_{o}\in\mathbb{R}^{|V|\times d}$ is the output weight matrix, $b_{o}\in\mathbb{R}^{|V|}$ is the bias vector, and $|V|$ is the size of the vocabulary. The next rule $\tau_{t}$ is then sampled from this distribution, $\tau_{t}\sim p_{\theta}(\tau_{t}|\tau_{1},\dots,\tau_{t-1})$, and fed back into the decoder as input for the next step, until the sequence is completed. After $H$ steps, the full sequence $\tau=(\tau_{1},\ldots,\tau_{H})$ is generated with probability $p_{\theta}(\tau)=\prod_{t=1}^{H}p_{\theta}(\tau_{t}|\tau_{1},\dots,\tau_{t-1})$, with the convention that $\tau_{1}$ is a special start symbol.
@@
Connection to equivalence-aware learning in DRL. Several techniques have been introduced to reduce the variance of policy gradient estimates. One widely used approach is the control variate method, where a baseline is subtracted from the reward to stabilize the gradient [DBLP:conf/uai/WeaverT01]. Another idea is Rao-Blackwellization [8a19a051-2be1-32f8-8b83-075ff1f85473]. Other approaches, such as reward reshaping [DBLP:conf/nips/ZhengOS18], modify rewards for specific state-action pairs. Inspired by stochastic variance-reduced gradient methods [johnson2013accelerating, DBLP:conf/iclr/0002FKLL21], [DBLP:conf/icml/PapiniBCPR18] proposed a variance-reduction technique tailored for policy gradients.
@@
B.4.3 Implementation of Egg-LLM
Our Egg is adopted on top of the original LLM-SR [DBLP:journals/iclr/llmsr25].
@@ B.4.3 Implementation of Egg-LLM
    Egg-based Feedback. Inferring Equivalent Expressions. For ease of integration with the e-graph system, we convert the generated Python functions into lambda expressions. This simplifies their manipulation and evaluation during equivalence analysis.
    1  # Original function format2  def hypothesis(x1, x2, coeffs):3   expr = x1 * coeffs[0] + x2 * coeffs[0]4   return expr5  6  # Converted into lambda format7  hypothesis = lambda x1, x2, coeffs: x1 * coeffs[0] + x2 * coeffs[0]This conversion allows us to transform the hypothesized Python function into an expression tree directly. For each fitted expression, we construct an e-graph and apply a predefined set of rewrite rules until reaching a fixed iteration limit. From the resulting e-graph, we sample $K$ unique equivalent expressions. Each sampled expression is then converted back into a Python function to facilitate further interaction with the LLM.
    1  # obtained equivalent expression2  equiv_seq = ["A->log(A)", "A->A*A", "A->x1", "A->x2"]3  # Converted function format4  def equiv_hypothesis(x1, x2, coeffs):5   return x1 * x2
    In subsequent iterations, the LLM receives feedback consisting of previously generated expressions along with their fitness scores. Our feedback is enriched with both the original hypotheses and their equivalent expressions derived via Egg module, enabling the model to refine future generations. High-performing expressions are retained and further updated over multiple rounds.
@@
B.4.4 Final Remark
The E2E-Transformer framework [DBLP:conf/nips/KamiennydLC22] is structurally very similar to the neural network used in the DRL model, with the main difference being that it is trained using a cross-entropy loss function. A straightforward way to integrate the Egg module into E2E-Transformer is through data augmentation, where additional training examples are generated using Egg.
To date, only one work has explored the use of a text-based diffusion model for symbolic regression [bastiani2025diffusion]. However, its implementation has not been released publicly. Moreover, the proposed pipeline relies on multiple interconnected components, which makes it difficult to reproduce or isolate for independent evaluation. For these reasons, we do not include a comparison between the text-based diffusion approach and Egg in this paper.
Several extensions of the baseline considered in this study offer empirical improvements but lack theoretical guarantees. For example, tenachi2023deep employ genetic programming to refine the predictions of the DRL model, and tenachi2023deep introduce unit constraints in physical systems to eliminate infeasible expressions. Integrating Egg with these variants is an interesting direction that we leave for future investigation.
Appendix C Experiment Settings
C.1 Baselines and Hyper-parameters Configurations
@@
MCTS. [DBLP:conf/iclr/Sun0W023] propose to use the Monte Carlo tree search algorithm to explore the space of symbolic expressions defined by a context-free grammar (described in Section 2) to find high-performing expressions. Despite being implemented similarly, this method appears under various names in previous works [DBLP:conf/icml/TodorovskiD97, DBLP:conf/dis/GanzertGSK10, DBLP:journals/kbs/BrenceTD21, DBLP:conf/iclr/Sun0W023, DBLP:conf/icml/KamiennyLLV23]. We choose the implementation  [^4] https://github.com/isds-neu/SymbolicPhysicsLearner from DBLP:conf/iclr/Sun0W023 and serve as a representative of this family of methods.
DRL. [DBLP:conf/iclr/PetersenLMSKK21] propose to use a recurrent neural network (RNN) trained via a (risk-seeking) reinforcement learning objective. The RNN sequentially generates candidate expressions and is optimized using a risk-seeking policy gradient to encourage the discovery of high-quality expressions. We chose this implementation  [^5] https://github.com/dso-org/deep-symbolic-optimization. We use three types of sequential decoders for the time benchmark setting. The major configurations are listed in Table 4.
@@
LLM-SR. [DBLP:journals/iclr/llmsr25] proposed to use pretrained large language models, such as GPT3.5, to generate symbolic expressions based on task-specific prompts. The expressions are evaluated and iteratively refined over multiple rounds. We choose their implementation at  [^6] https://github.com/deep-symbolic-mathematics/LLM-SR.
@@
Expression-related Configurations. When fitting the values of open constants in each expression, we sample a batch of data with batch size $1024$ from the data Oracle. The open constants in the expressions are fitted on the data using the BFGS optimizer [^7] https://docs.scipy.org/doc/scipy/reference/optimize.minimize-bfgs.html. We use a multi-processor library to fit multiple expressions using 8 CPU cores in parallel. This greatly reduced the total training time.
@@ C.2 Evaluation Metrics
where the empirical variance $\sigma_{y}=\sqrt{\frac{1}{n}\sum_{i=1}^{n}\left(y_{i}-\frac{1}{n}\sum_{i=1}^{n}y_{i}\right)^{2}}$. Note that the coefficient of determination ($R^{2}$) metric [nagelkerke1991note, DBLP:conf/nips/CavaOBFVJKM21] is equal to $(1-\text{NMSE})$ and therefore omitted in the experiments.
Appendix D Extra Experimental Results
D.1 Visualization of Egg Construction
@@ I.44.4 $c_{1}x_{0}x_{1}\log(x_{2}/x_{3})$ $c_{1}x_{0}x_{1}(\log(x_{2})-\log(x_{3}))$ Figure 18
D.2 Additional Visualization for Selected Expressions in Feynman Dataset